# Variations on quantum de Finetti theorems and operator valued Martin boundaries: a Choquet–Deny approach

### v2

@@ A.1. A generalized quantum de Finetti theorem for general $C^{*}$-algebras
If $b\in\mathcal{B}_{+}$ is invertible, then $K_{\mathcal{A}^{*},b}$ is the closed convex hull of $\mathrm{ex}(K_{\mathcal{A}^{*},b})$.
@@
Using as in Section 3, the results of Chapter 13 of [11], we get the following representation theorem:
@@
Every $\chi\in K_{\mathcal{A}^{*},b}\backslash\{0\}$ can be represented by a maximal Borel probability measure on $\mathrm{ex}(K_{\mathcal{A}^{*},b})$ (in the sense of Pro. 3.1).
@@
$\mathrm{ex}(K_{\mathcal{A}^{*},b})$ is contained in $\{(\chi\otimes\beta^{\otimes l})_{l=0}^{\infty}\mid\chi\in\mathcal{A}^{*}_{+},\beta\in\mathcal{B}^{*}_{+}\}$.
@@ A.2. Infinite-dimensional Biane–Choquet–Deny theorem
We say that $x\in\mathcal{A}_{*}\otimes\mathcal{U}(G)$ is positive if $(\mathrm{id}_{\mathcal{A}^{*}}\otimes\pi)(x)$ is a positive linear form in $\mathcal{A}^{*}\otimes B(\mathcal{H})=(\mathcal{A}\otimes B(\mathcal{H}))^{*}$ for any finite-dimensional unitary representation $(\pi,\mathcal{H})$ of $G$. Then, by extending the proof of Proposition 4.1, we can check that $P_{\mathcal{A}^{*},\rho}$ is a positive map.
@@
$B_{\mathcal{A}^{*},\rho}$ is the closed convex hull of $\mathrm{ex}(B_{\mathcal{A}^{*},\rho})$. Moreover, every $x\in B_{A^{*},\rho}\backslash\{0\}$ is represented by a maximal Borel probability measure on $\mathrm{ex}(B_{\mathcal{A}^{*},\rho})$ (in the sense of Prop. 3.1).
@@
The extreme points $\mathrm{ex}(B_{\mathcal{A}^{*},\rho})$ are contained in $\mathcal{A}^{*}_{+}\otimes\mathcal{E}_{\rho}(G)$. In particular, $B_{\mathcal{A}^{*},\rho}$ is the closed convex hull of $\mathcal{A}^{*}_{+}\otimes\mathcal{E}_{\rho}(G)$.

### v3

@@ A.1. A generalized quantum de Finetti theorem for general $C^{*}$-algebras
If $b\in\mathcal{B}_{+}$ is invertible, then $K_{\mathcal{A}^{*},b}$ is the closed convex hull of $\mathrm{exr}(K_{\mathcal{A}^{*},b})$.
@@
By the same proof of Proposition 3.1, we obtain the following representation theorem:
@@
Every $\chi\in K_{\mathcal{A}^{*},b}\backslash\{0\}$ can be represented by a Borel probability measure $M_{\chi}$ on $K_{\mathcal{A}^{*},b}$ such that $M_{\chi}$ is supported on any closed subset of $K_{\mathcal{A}^{*},b}$ which contains $\mathrm{exr}(K_{\mathcal{A}^{*},b})$.
@@
$\mathrm{exr}(K_{\mathcal{A}^{*},b})$ is contained in $\{(\chi\otimes\beta^{\otimes l})_{l=0}^{\infty}\mid\chi\in\mathcal{A}^{*}_{+},\beta\in\mathcal{B}^{*}_{+}\}$.
@@ A.2. Infinite-dimensional Biane–Choquet–Deny theorem
We say that $x\in\mathcal{A}^{*}\otimes\mathcal{U}(G)$ is positive if $(\mathrm{id}_{\mathcal{A}^{*}}\otimes\pi)(x)$ is a positive linear form in $\mathcal{A}^{*}\otimes B(\mathcal{H})=(\mathcal{A}\otimes B(\mathcal{H}))^{*}$ for any finite-dimensional unitary representation $(\pi,\mathcal{H})$ of $G$. Then, by extending the proof of Proposition 4.1, we can check that $P_{\mathcal{A}^{*},\rho}$ is a positive map.
@@
$B_{\mathcal{A}^{*},\rho}$ is the closed convex hull of $\mathrm{exr}(B_{\mathcal{A}^{*},\rho})$. Moreover, every $x\in B_{A^{*},\rho}\backslash\{0\}$ is represented by a Borel probability measure on $\mathrm{exr}(B_{\mathcal{A}^{*},\rho})$ (in the sense of Prop. 3.1).
@@
The extreme points $\mathrm{exr}(B_{\mathcal{A}^{*},\rho})$ are contained in $\mathcal{A}^{*}_{+}\otimes\mathcal{E}_{\rho}(G)$. In particular, $B_{\mathcal{A}^{*},\rho}$ is the closed convex hull of $\mathcal{A}^{*}_{+}\otimes\mathcal{E}_{\rho}(G)$.